OpenStudy (anonymous):

A motorboat took 3 hours to make a downstream trip with a current of 6 km/hr. The return trip against the same current took 5 hours. Find the speed of the boat in still water.

OpenStudy (anonymous):

I got 24, but I don't know if its right

OpenStudy (anonymous):

r*t=d Downstream (r+6)(3) Upstream (r-6)(5) The distance is the same (r+6)(3)=(r-6)(5) 3r+18=5r-30 48=2r r=24 The boat is traveling at 24 km/hr

OpenStudy (anonymous):

ok thank you very much :)